OpenStudy (anonymous):

find and simlify each of the following, given f(x)=sqrtx+5 and g(x)=5x+1 question 1 f^-1(x) question 2 (g0f)(-1) that is the little circle between g and f

OpenStudy (blockcolder):

\[f(x)=\sqrt{x}+5\\ y=\sqrt{x}+5\\ x=\sqrt{y}+5\\ x-5=\sqrt{y}\\ y=f^{-1}(x)=(x-5)^2\]

OpenStudy (blockcolder):

\[(g \circ f)(-1)=g(f(-1))\] However, f(-1) is undefined, so g(f(-1)) is also undefined.

OpenStudy (anonymous):

dont you put 5x+1 for g and multiply it to sqrt x+5 then all that by -1

OpenStudy (anonymous):

so the answer to both is undefined

OpenStudy (anonymous):

thank you blockcolder